# LiveMCPBench: Can Agents Navigate an Ocean of MCP Tools?

### v1

@@ Introduction
Table 3: Performance efficiency metrics: Steps denotes average dialogue turns, Tools indicates average used tools, execute represents average tool executions, route refers to average retrievals, and Overall shows average task success rate.
Figure 5: Log-Price vs. Performance scatter plot with pareto frontier representation. Different colors represent different model families.
Efficiency Analysis
@@
To compare the behavioral characteristics of different models, we present the average number of dialogue turns, used tools, tool execution attempts, and retrieval calls in Table 3. Based on these metrics, we draw the following conclusions:
  1. 1.
    Claude series models exhibit more proactive exploration and utilization behavior. Their retrieval and execution frequencies are significantly higher than other models, accompanied by a greater number of used tools. This suggests that Claude models actively engage with and adapt to the tool-augmented environment, demonstrating a stronger tendency to explore and exploit available tools.
  2. 2.
    Most models suffer from severe underutilization of tools. The average number of tools used by these models remains close to 1, indicating that once a model identifies and adopts a single tool, it tends to rely exclusively on it while neglecting other available tools. This behavior highlights a critical limitation in their ability to dynamically leverage multiple tools during task execution.
@@
In practical applications, a trade-off between model performance and cost must be carefully considered. To provide actionable insights for model selection, we plotted the relationship between logarithmic cost and performance, along with the corresponding pareto frontier (Lotov, Bushenkov, and Kamenev 2004). As illustrated in Figure 5, our analysis reveals two key findings:
  1. 1.
    Near-Linear Trade-off on the Pareto Frontier. The performance and logarithmic cost of models along the pareto frontier exhibit an approximately linear relationship. This observation presents a valuable opportunity for optimizing cost-performance balance in real-world tool-calling agent.
  2. 2.
    Optimal Cost-Performance Models. The models positioned on the pareto frontier represent the most cost-effective choices for tool calling. These include Qwen3-32B, Qwen2.5-72B-Instruct, Deepseek-R1-0528, and Claude-Sonnet-4, each demonstrating distinct advantages in terms of cost-performance efficiency.
Error Analysis
Figure 6: Distribution of errors for four types. We counted two agent trajectories for the advanced claude series models.
@@
We conducted a detailed error analysis on the trajectories of current retrieval and invocation agents to provide insights for future development. Human annotators were employed to classify error types in the trajectories of Claude-Opus-4 and Claude-Sonnet-4. Based on the modules defined in the MCP Copilot Agent framework, we identified four distinct and easily distinguishable error categories (Figure 6). Each erroneous trajectory was uniquely classified into one error type without overlap. Detailed error examples are provided in Appendix G.
@@
Retrieve errors arise when semantically appropriate queries fail to match available tools due to retrieval system shortcomings. For example, in the task “Convert the YouTube video to MP3 format,” the retrieval system may overlook the youtube downloader tool (which supports format conversion) due to unrecognized semantic equivalence between “convert to MP3” and the tool’s documented “extract audio tracks” functionality. These errors highlight challenges in hierarchical retrieval (e.g., MCP server-tool structures) and semantic similarity computation. Dominating the error distribution, retrieve errors underscore the critical need for enhanced retrieval architectures and more robust similarity metrics.
@@
Related Work
Tool-Use Benchmarks
Most existing benchmarks for tool utilization rely on simulated APIs due to the inherent instability of real-world APIs. For instance, API-Bank (Li et al. 2023) and StableToolBench (Guo et al. 2024) employ artificially constructed toolsets to ensure API stability. Other works, such as ToolAlpaca (Tang et al. 2023) and Seal-Tools (Wu et al. 2025), collect real-world API interfaces but are unable to execute actual calls. A third category of tool-use benchmarks, including ToolBench (Qin et al. 2024) and ShortcutsBench (SHEN et al. 2025), attempts to integrate real-world APIs but faces challenges due to rapid API changes, leading to frequent tool unavailability (Guo et al. 2024).
Recent efforts, like StableToolBench-MirrorAPI (Guo et al. 2025), leverage fine-tuned LLMs to simulate API interfaces and calls. However, these prior tool-use benchmarks predominantly focus on API-based tools, which introduces instability and limits functionality—particularly in directly manipulating local user files or enabling complex operations (e.g., interacting with local software).
The emergence of MCP has shifted this paradigm by providing a stable, unified interface, enabling the development of general-purpose toolsets. In this work, we construct a practical MCP toolset, addressing both the instability of APIs and their functional constraints, thereby delivering a comprehensive and reliable real-world tool-use benchmark.
MCP Benchmarks
The evaluation of MCP systems remains an emerging and rapidly evolving field. Among existing benchmarks, MCPBench (Luo et al. 2025) stands as one of the earliest efforts, primarily focusing on comparative analyses between MCP tools and traditional API-based tools. Building upon this, MCP-RADAR (Gao et al. 2025) extends the evaluation scope by introducing a multi-dimensional assessment framework that examines critical aspects such as efficiency, accuracy, and robustness. More recently, MCPEval (Liu et al. 2025) has been advanced the field further by proposing a fine-grained evaluation framework capable of automatically generating queries to assess the performance of MCP servers.
Despite these advancements, existing benchmarks suffer from a key limitation: their evaluations are predominantly conducted on small-scale MCP servers (typically around 10 servers), which inadequately reflects real-world scenarios where agents must operate in large-scale, dynamic environments. To bridge this gap, our work introduces a large-scale MCP toolset and systematically investigates agent capabilities in accomplishing everyday tasks through extensive tool utilization.
Recent efforts, such as RAG-MCP (Gan and Sun 2025), MCPZero (Fei, Zheng, and Feng 2025), and ScaleMCP (Lumer et al. 2025), have explored retrieval methods over large-scale MCP toolsets. However, these approaches are constrained by rigid pipelines that lack dynamic adaptability in tool invocation and error feedback. Furthermore, ScaleMCP relies on a manually constructed toolset with limited functional diversity, restricting its applicability to broader, real-world use cases.
Conclusion
In this paper, we present LiveMCPBench, a benchmark designed to evaluate the capability of agents in accomplishing daily tasks using large-scale MCP toolset. We introduce LiveMCPTool, a comprehensive and readily deployable collection of MCP tools. We propose LiveMCPEval, an automated evaluation framework based on the LLM-as-a-Judge, which effectively assesses complex tool-usage tasks characterized by time-varying dynamics and diverse completion paths. Human evaluations confirm the reliability of LiveMCPEval. Furthermore, we develop MCP Copilot Agent, an agent framework capable of autonomous exploration and dynamic decision-making in large-scale MCP environments. We conduct extensive evaluations on ten frontier models, revealing limitations in widely-used LLMs when applied to large-scale tool invocation tasks. Our in-depth analysis uncovers distinct behavioral patterns across different models and identifies the most cost-effective solutions. Finally, through detailed error analysis, we highlight two critical shortcomings in current models: (1) deficiencies in task decomposition and planning, and (2) inadequate adaptation of tool retrieval systems in MCP environments. These findings provide clear directions for future improvements in the field.
@@ Appendix A Limitations
The LiveMCPEval component relies heavily on LLM-based assessment. Although we have validated the evaluation accuracy through human experiments, potential model biases may still influence the results. To mitigate this concern, we conducted extensive case studies analyzing model judgment failure cases, which helps improve the robustness of our evaluation framework.
@@ Appendix B Ethical Considerations
Appendix C Details of Task Construction
C.1 Task Statistics
Figure 7: Task distribution in LiveMCPBench: a comprehensive benchmark comprising 95 tasks Across 6 distinct domains.
@@ C.1 Task Statistics
C.2 Annotation Principles
LiveMCPBench focuses on leveraging a large-scale MCP toolset to accomplish complex tasks. To ensure high-quality task construction, we employ two groups of annotators involving proposers and verifiers. All annotators first freely explore the MCP toolset, including tool descriptions and real-world calls, to gain familiarity with its functionalities.

### v2

@@ 3.1. Setup
Table 2. Performance and efficiency metrics: Steps denotes average dialogue turns, Tools indicates average used tools, execute represents average tool executions, route refers to average retrievals, Tokens refers to total tokens consumed (in millions), and Overall shows average task success rate.
Model Steps Tools excute route Tokens (M) Overall (%)
Claude-Sonnet-4-20250514 20.09 2.71 5.59 2.98 6.07 78.95
Claude-Opus-4-20250514 25.53 3.40 6.93 4.35 7.16 70.53
GPT-5 20.89 1.77 7.24 2.91 6.84 52.63
Qwen3-235B-A22B 10.33 1.24 2.11 2.00 2.88 48.42
DeepSeek-R1-0528 16.76 1.59 5.12 1.77 3.73 48.42
GPT-5-Mini 11.18 1.28 2.88 1.37 3.63 45.26
GPT-4.1-Mini 10.89 1.37 2.71 1.65 2.96 44.21
Qwen2.5-72B-Instruct 11.22 1.31 2.80 1.38 3.55 43.16
DeepSeek-V3-0324 8.33 1.01 1.29 1.41 2.11 42.11
Gemini-2.5-Pro 8.08 0.99 1.46 1.35 2.03 41.05
GPT-4.1 9.03 1.31 1.72 1.64 2.48 38.95
Qwen3-32B 9.99 1.16 2.31 1.19 1.93 30.53
We evaluate 12 frontier models: Claude-Opus-4 and Claude-Sonnet-4 (Anthropic, 2025), GPT-4.1 and GPT-4.1-Mini (Openai, 2025a), GPT-5 and GPT-5-Mini (Openai, 2025b), Gemini-2.5-Pro (Google, 2025), Deepseek-V3 and Deepseek-R1 (DeepSeek-AI, 2025), Qwen3-235B-A22B and Qwen3-32B (Yang et al., 2025), and Qwen2.5-72B-Instruct (Qwen et al., 2025). For evaluation, we employ Deepseek-V3 as default evaluator.
Tool Retrieval Configuration. To ensure consistent retrieval performance across experiments, we adopt a standardized configuration. Specifically, we employ Qwen2.5-72B-Instruct to generate server-level tool descriptions and Qwen3-Embedding-0.6B (Zhang et al., 2025) to obtain dense embeddings. In all experiments, the retrieval system returns the top-$5$ tools most relevant to the query, ensuring a controlled and comparable evaluation setup.
3.2. Main Results
@@
To compare the behavioral characteristics of different models, we present the average number of dialogue turns, used tools, tool execution attempts, retrieval calls and task success rates in Table 2. Based on these metrics, we draw the following conclusions:
  1. (1)
    Performance Variance Across Models. We observe substantial variance in performance: while most contemporary models achieve only 30%–50% task success rates, the Claude series substantially outperforms them.
  2. (2)
    Active tool composition correlates with stronger performance. Claude-Sonnet-4, with 2.71 tools, 5.59 executions, and 2.98 retrievals, achieves the highest success rate, far outperforming conservative models such as Gemini-2.5-Pro or DeepSeek-V3, which use only one tool on average. This suggests that proactive interaction with the tool environment is a key mechanism for effective MCP usage.
  3. (3)
    Efficiency hinges on balancing exploration and exploitation. While Claude-Opus-4 explores even more aggressively (3.40 tools), its success rate drops to 70.53%, indicating diminishing returns and potential error accumulation. Conversely, models that under-explore (e.g., Qwen3-32B at 1.16 tools) fail to adapt. Thus, performance is not determined by raw tool counts but by how effectively a model manages the exploration-exploitation trade-off.
Figure 3. Correlation between evaluation performance and human agreement across evaluators on Claude trajectories.
Is LiveMCPEval Reliable? We evaluated the reliability of LiveMCPEval by comparing its automatic judgments with human annotations on execution trajectories of top Claude models and by testing models as evaluators (Figure 3). Results show that LiveMCPEval is accurate under appropriate evaluators: Deepseek-V3 reaches 78.95% human agreement, while GPT-4.1 Mini and Qwen2.5-72B-Instruct achieve $\sim$75%. In contrast, Deepseek-R1, Claude-Opus-4, and Qwen3-32B yield lower agreement (60–70%), likely due to difficulty handling long trajectories. Overall, LiveMCPEval is dependable, but its effectiveness depends on the model.
Is LiveMCPEval Generalizable? We tested the generalizability of LiveMCPEval on Claude-Sonnet-4’s trajectories by examining how LLM-generated key points affect human agreement (Figure 4, examples in Appendix H). Results show that LiveMCPEval generalizes well across models: even without human-annotated references, most evaluators achieved higher agreement using automatically generated key points. This demonstrates the framework’s openness to new tasks and tools, since auto-generated key points can adaptively serve as substitutes when human references are unavailable. Moreover, Deepseek-V3 best leverages human-annotated key points, suggesting advantages for scenarios where such references are available.
4. Analysis
4.1. Efficiency Analysis
Figure 4. Human agreement rates of different evaluators when using manual versus LLM-generated key points.
In practical applications, a trade-off between model performance and cost must be carefully considered.
Table 2 reports the average number of steps and total token consumption for each model. As expected, token usage increases with the steps. However, we observe that Claude models and GPT-5 consume substantially more tokens (6–7M) compared to the other models (2–3M). Beyond their higher average step counts, their more aggressive tool-use behavior also contributes to the increased token usage. For example, Claude-Sonnet-4 requires an average of 20.09 steps and consumes 6.70M tokens, whereas Qwen3-235B requires 16.76 steps and consumes only 3.73M tokens. This large gap is largely explained by their tool-use patterns: Claude-Sonnet-4 invokes an average of 2.71 tools per task, while Qwen3-235B uses only 1.59. These findings suggest that more aggressive tool exploration leads to substantially higher token consumption, due to factors such as more exploratory trials and error-recovery attempts.
@@ 4.1. Efficiency Analysis
To provide actionable insights for model selection, we plotted the relationship between logarithmic cost and performance, along with the corresponding Pareto frontier (Lotov et al., 2004). As illustrated in Figure 5. The models positioned on the Pareto frontier represent the most cost-effective choices. These models demonstrate distinct advantages in terms of cost-performance efficiency.
4.2. Impact of Tool Retrieval Configuration
To assess the impact of retriever configurations on the agent, we conduct ablation experiments on two factors: the number of retrieved tools and the choice of embedding model. Results are summarized in Table 3. For all paired comparisons against the main setting, we apply an exact McNemar test (McNemar, 1947) on per-instance success outcomes, which is appropriate for our matched binary predictions and avoids asymptotic assumptions. The resulting $p$-values provide evidence for the robustness of our primary hyperparameter choice.
Figure 5. Log-Price vs. Performance scatter plot with Pareto frontier. Colors represent model families.
Regarding $k$, reducing the number of retrieved tools from our default setting of $5$ to $1$ leads to a performance degradation, with the success rate dropping from $78.95\%$ to $64.21\%$. This decline is statistically significant under the exact McNemar test ($p=0.02<0.05$). In contrast, increasing $k$ further to $10$ yields no additional gain, as the success rate plateaus at $78.95\%$ ($p=1.00$). This suggests that our primary configuration is already near-optimal. We attribute this saturation to the retriever’s limited ability to surface additional feasible tools, rather than a lack of candidates in the pool.
For the embedding model, substituting Qwen3-Embedding-0.6B with BGE-M3 (Chen et al., 2025) leads to only a minor decrease in performance (from $78.95\%$ to $76.84\%$), and the difference is not significant by the exact McNemar test ($p=0.84$). Taken together, these statistically grounded comparisons suggest that our main experimental setting is stable to reasonable variations in both $k$ and embedding choice, and further support the conclusion that the dominant bottleneck lies in the current tool-retrieval methodology itself rather than in a specific embedding model or hyperparameter.
4.3. Single Evaluator Bias Analysis
Although the reliability of LiveMCPEval has been preliminarily validated through human agreement rates, further in-depth analysis of its applicability and potential improvements is crucial for the long-term sustainability of the evaluation process. The potential impact of evaluator bias on the reliability of LLM-as-Judge has raised concerns about the robustness of such evaluation systems.
A potential solution is majority voting. To assess the effectiveness of this approach, we used all 10 models from the main experiment as evaluators to assess the trajectory of Claude-Sonnet-4. The default evaluator, deepseek-v3, has a human agreement rate of 81.05%, while the weakest evaluator, deepseek-r1, has a human agreement rate of 58.95%. We considered results with at least a half vote as predictions, yielding a human agreement rate of 77.89%. This rate is slightly lower than that of the default evaluator but significantly higher than the weakest evaluator. To further explore the upper performance limit of multiple models, we calculated the maximum human agreement rate (where any model output agreeing with human results is considered correct) to be 97.89%. These findings suggest that when the performance of individual evaluators is uncertain, using a multi-model voting approach is a more robust choice.
To further verify whether the specific model bias of our default deepseek-v3 evaluator affects the relative ranking of performance, we also used Qwen2.5-72B-Instruct as an evaluator and reassessed all trajectories. The experimental results, shown in Table 4, indicate that the relative rankings remained almost unchanged.
To quantitatively examine the consistency between the two evaluator, we further computed Kendall’s $\tau$-b (Kendall, 1948), which measures the ordinal association between two ranked lists while accounting for ties. The result ($\tau$-b = 0.8837 p = 0.00008) indicates a strong and statistically significant agreement between the rankings. In other words, although individual evaluators may exhibit model-specific preferences, their induced orderings over trajectories are largely aligned.
Table 3. Ablation study on the number of retrieved tools ($k$) and embedding model selection. We employ Claude-Sonnet-4 as the agent and DeepSeek-V3 as the evaluator.
Settings Overall (%) McNemar p-value
Qwen3-Embedding-0.6B + kk=5 (main) 78.95 -
Change $k$
Qwen3-Embedding-0.6B + kk=1 64.21 0.02
Qwen3-Embedding-0.6B + kk=10 78.95 1.00
Change Embedding Model
BGE-M3 + kk=5 76.84 0.84
4.4. Evaluation Error Analysis
Figure 6 presents the confusion matrices between the model-based evaluator (DeepSeek-v3) and human annotators for the Claude family of models. The evaluation of claude-sonnet-4 achieved an F1 score of 87.32, while claude-opus-4 reached 83.08, underscoring the relatively high reliability of model-based evaluation. To better understand the observed discrepancies between human judgments and automatic evaluation, we conducted a detailed error analysis.
For clarity and mutual exclusiveness, we categorize evaluation errors into the following four types:
  • Output Completeness Error (30%): Cases where the evaluator disagreed with humans on whether the model output provided a fully correct and final solution, versus only partial or intermediate steps.
  • Hallucinated Completion (45%): Cases where the agent incorrectly assumed completion of a task or access to unavailable information, leading the evaluator to mistakenly align with this hallucination.
  • Granularity Mismatch (17.5%): Cases where the evaluator and humans diverged due to differences in tolerance for detail, such as penalizing superficial formatting inconsistencies or overlooking missing justifications.
  • Others (7.5%): Residual errors, primarily attributable to human annotation mistakes rather than systematic evaluator weaknesses.
Table 4. Reevaluation of main experiment results using Qwen2.5-72B-Instruct. The relative rankings exhibit strong consistency with the original evaluation (Kendall’s $\tau$-b = 0.8837, $p<0.001$).
Model Overall (%) Origin Rank New Rank
Claude-Sonnet-4 77.89 1 1
Claude-Opus-4 64.21 2 2
GPT-5 51.58 3 3
Qwen3-235B 40.00 4 4
GPT-5-Mini 40.00 6 5
Qwen2.5-72B 38.95 8 6
Deepseek-R1 37.89 5 7
GPT-4.1-Mini 31.58 7 8
Deepseek-V3 31.58 9 9
Gemini-2.5-Pro 29.79 10 10
GPT-4.1 27.37 11 11
Qwen3-32B 25.26 12 12
Figure 6. Confusion matrices comparing model-based evaluation (DeepSeek-v3) against human annotators.
4.5. Error Analysis
@@
We conducted a detailed error analysis on the trajectories of current retrieval and invocation agents to provide insights for future development. Human annotators were employed to classify error types in the trajectories of Claude-Opus-4 and Claude-Sonnet-4. Based on the modules in the MCP Copilot Agent framework, we identified four distinct error categories (Figure 7). Each erroneous trajectory was uniquely classified into one error type without overlap. Detailed error examples are provided in Appendix I.
@@
Retrieve errors arise when semantically appropriate queries fail to match available tools due to retrieval system shortcomings. For example, in the task “Convert the YouTube video to MP3 format,” the retrieval system may overlook the youtube downloader tool (which supports format conversion) due to unrecognized semantic equivalence between “convert to MP3” and the tool’s documented “extract audio tracks” functionality. These errors highlight challenges in hierarchical retrieval (e.g., MCP server-tool structures) and semantic similarity computation. Dominating the error distribution, retrieval errors underscore the critical need for enhanced retrieval architectures and more robust similarity metrics.
@@
Figure 7. Distribution of errors for four types. We counted trajectories of Claude series models.
5. Related Work
5.1. Tool-Use Benchmarks
Most existing benchmarks for tool use rely on simulated APIs, given the instability of real-world interfaces. For example, API-Bank (Li et al., 2023) and $\tau$-bench (Yao et al., 2025) construct artificial toolsets to ensure stability, while ToolAlpaca (Tang et al., 2023) and Seal-Tools (Wu et al., 2025) collect real-world APIs but cannot execute actual calls. A third line of work, such as ToolBench (Qin et al., 2024) and ShortcutsBench (SHEN et al., 2025), attempts to integrate real APIs, but frequent interface changes often render tools unusable (Guo et al., 2024). More recent efforts, e.g., StableToolBench-MirrorAPI (Guo et al., 2025), use fine-tuned LLMs to simulate APIs and calls. Yet, these benchmarks remain largely API-centric, inheriting instability and lacking support for broader functionality—such as direct manipulation of local files. The emergence of MCP offers a new paradigm: a stable, unified interface for building general-purpose toolsets. In this work, we leverage MCP to construct a practical toolset that overcomes both API instability and functional limitations, enabling a comprehensive and reliable real-world tool-use benchmark.
5.2. MCP Benchmarks
The evaluation of MCP systems is still in its early stages and continues to evolve rapidly. Among existing efforts, MCPBench (Luo et al., 2025) represents one of the first benchmarks, mainly comparing MCP tools with traditional API-based tools. Building on this, MCP-RADAR (Gao et al., 2025) broadens the scope by introducing a multi-dimensional framework that evaluates efficiency, accuracy, and robustness. More recently, MCPEval (Liu et al., 2025) has advanced the field with a fine-grained framework capable of automatically generating queries to assess MCP server performance. Despite these advances, current benchmarks share a key limitation: they typically evaluate only small-scale MCP servers (around 10 servers), which fail to capture real-world settings where agents must operate in large, dynamic MCP ecosystems. To address this gap, we introduce a large-scale MCP toolset and systematically examine agent capabilities in accomplishing daily tasks through tool use. Recent work, such as RAG-MCP (Gan and Sun, 2025), MCPZero (Fei et al., 2025), and ScaleMCP (Lumer et al., 2025) has begun exploring retrieval over large-scale MCP toolsets. However, these approaches are constrained by rigid pipelines that lack adaptability in tool invocation and error recovery. Moreover, ScaleMCP depends on a manually constructed toolset with limited functional diversity, reducing its applicability to broader real-world scenarios.
6. Conclusion
We presented LiveMCPBench, a benchmark consisting of 95 daily tasks for evaluating tool-use agents in large-scale MCP ecosystems. Alongside, we released LiveMCPTool, a ready-to-use collection of 527 tools, and proposed LiveMCPEval, an automatic evaluation framework that handles task dynamism and solution diversity. We further built the MCP Copilot Agent to study tool retrieval and composition in large-scale MCP settings. Our experiments across 12 frontier models highlight persistent limitations: tool retrieval remains the primary bottleneck, while effective tool composition proves equally critical. These findings underscore the need for future systems that jointly advance tool retrieval and foster robust tool composition capabilities.
@@ Appendix A Limitations
The LiveMCPEval component relies heavily on LLM-based evaluation. Although we have validated the accuracy through human experiments, potential model biases may still influence the results. To mitigate this concern, we conducted case studies analyzing model judgment failure cases, which helps improve the robustness of our evaluation framework.
@@
Server Reliability and Maintenance.
While our framework highlights the “plug-and-play” nature of LiveMcpTool, we acknowledge that the current work does not explicitly address long-term aspects such as security, versioning, and maintenance of deployed servers. In practice, sustainable applicability requires systematic guidelines for server updates, monitoring, and health curation. To ensure reproducibility of our reported results, we provide a fixed and validated toolset version that has undergone security checks, which guarantees a safe and consistent experimental environment. Beyond this controlled release, we view long-term sustainability as an orthogonal but essential research direction: future iterations of LiveMCPBench could incorporate auditing modules or best-practice recommendations (e.g., periodic health checks, semantic versioning, and community-curated registries) to further ensure secure and reliable deployment.
@@ Appendix B Ethical Considerations
Appendix C LLM Usage
We used LLMs, specifically OpenAI’s GPT-based tools, only to check grammar, improve readability, and polish draft wording. All conceptual, technical, and scientific contributions are entirely the authors’ own.
Appendix D Reproducibility statement
We open-source all the code and human annotations used in our experiments. In addition, we include the complete set of tasks and the entire LiveMCPTool. Furthermore, Appendix G details the computational resources we utilized, the access methods for private model APIs.
Appendix E Details of Task Construction
E.1. Task Statistics
Figure 8. Task distribution in LiveMCPBench: a benchmark comprising 95 daily tasks across 6 distinct domains.
Figure 9. Example of the task annotation process.
@@ E.1. Task Statistics
Importantly, the retained 95 tasks were selected not for sheer quantity, but for domain coverage, generality, and compositional complexity. For instance, beyond single-step retrieval, many tasks involve multi-hop reasoning (e.g., analyzing market data followed by producing a spreadsheet report) or cross-domain integration (e.g., combining travel planning with lifestyle queries). Compared to prior agent benchmarks, LiveMCPBench thus prioritizes both breadth of domains and depth of task composition, providing a balanced yet realistic testbed despite the smaller absolute number of tasks relative to very large-scale static datasets.
Annotator MCP Usage.
To accomplish the benchmark tasks, annotators utilized 55 out of the 70 available servers (78.57%) and 150 out of the 527 available tools (28.46%). This indicates that while not all resources were required, a substantial and diverse subset of servers and tools was actively engaged, reflecting realistic large-scale MCP tool usage.
In practice, many MCP servers contain noisy or redundant tools that are not essential for task completion, and our benchmark naturally captures this characteristic by relying on only the most relevant subset.
Figure 10. Distribution of servers in LiveMCPTool, categorized into 8 distinct types (70 servers).
Figure 11. Distribution of tools in LiveMCPTool, categorized into 8 distinct types (527 tools).
E.2. Annotation Principles
LiveMCPBench focuses on leveraging a large-scale MCP toolset to accomplish complex tasks. To ensure high-quality task construction, we employ two groups of annotators involving proposers and verifiers (see Figure 9). All annotators first freely explore the MCP toolset, including tool descriptions and real-world calls, to gain familiarity with its functionalities.